OpenStudy (anonymous):

Nadia took two practice exams to prepare for her college entrance exam. On the first practice exam, she scored 1,100. After studying, her score increased by 25%. What was her score on the second practice exam? a] 1,125 b] 1,350 c] 1,375 d] 1,467

OpenStudy (a_clan):

c

OpenStudy (anonymous):

= 1100 * 1.25 = 1375

OpenStudy (salina):

1100*0.25=275 275+1100=1375 is c

OpenStudy (anonymous):

thank you both so much (:!